OpenStudy (anonymous):

1. What percentage of the data falls outside 1 standard deviation of the mean?

OpenStudy (tkhunny):

Well, that would depend on the data being approximately Normal. Is it? Think "Empirical Rule".

OpenStudy (anonymous):

is it 1%

OpenStudy (tkhunny):

The Empirical Rule suggests that 68% of the data are WITHIN 1 Standard Deviation of the Mean.

OpenStudy (anonymous):

What percentage of the data falls outside 1 standard deviation of the mean?

OpenStudy (anonymous):

One standard deviation for one side will be 34% of data. So within 1 std. dev. to both sides will be 68% (approximately) .the data falls outside 1 standard deviation of the mean will be 1.00 - 0.68 = 0.32 (32 %)

OpenStudy (anonymous):

 This is assuming the empirical rule: 100-68=32